# A Two-Sink Solution to the Self-Similar Euler Equations

### v1

@@ Abstract
We construct the first known example of a self-similar solution to the two-dimensional incompressible Euler equations whose pseudo-velocity has more than one stagnation point. The solution is also a homogeneous steady state of the Euler equations. In contrast, any homogeneous steady state with bounded vorticity necessarily admits only a single stagnation point at the origin. Our construction develops cusps in the velocity along two lines passing through the origin, thereby allowing stagnation points other than the origin.
@@ 1 Introduction
$\begin{gathered}\partial_{t}u+(u\cdot\nabla)u+\nabla p=0\\ u(\cdot,t=0)=u_{0}(\cdot).\end{gathered}$ (1)
One solves for the velocity vector field $u(x,t)$ and scalar pressure field $p(x,t)$ starting from initial data $u_{0}(x)$. The incompressible Euler equations in two-dimensions model the motion of inviscid fluids in thin plates. Due to a transport structure of the equations, the vorticity, $\textrm{curl }u$, is transported by the velocity $u$, which facilitates the existence theory for weak solutions starting from quite irregular initial data (for example, the existence for $L^{1}\cap L^{p}$ vorticity was proven in [8]). On the other hand, uniqueness of solutions is only known when the vorticity is bounded, and non-uniqueness is only known (by convex integration arguments) due to a lack of energy conservation for $C^{1/3^{-}}$ velocity. More recently, the authors of [6] used convex integration to prove non-uniqueness for $L^{1+}$ vorticity data on $\mathbb{T}^{2}$. The question whether unbounded and integrable ($L^{1}\cap L^{p}$) vorticity data guarantees the uniqueness of weak solutions is open and referred to as the “Yudovich Problem”.
@@
$\begin{array}[]{rcl}(\nabla^{\perp}\Psi-\frac{1}{\alpha}\xi)\cdot\nabla_{\xi}\Omega&=&\Omega,\\ \Delta\Psi&=&\Omega.\end{array}$ (2)
@@
$\omega(x,t):=t^{-1}\Omega(xt^{-1/\alpha}),$ (3)
@@
Bifurcation in the self-similar variables is believed to be a plausible non-uniqueness mechanism for the Euler equations, especially since Vishik proved the non-uniqueness of solutions with time-integrable forcing in [14] (a result further refined by the authors of [1]). We now describe the bifurcation behavior necessary for non-unique solutions to arise. If one can construct two different self-similar profiles with the same “data at infinity” in the self-similar variables, then one has proven the existence of two different self-similar solutions to the Euler equations from the same homogeneous data.
@@
We have argued the importance of self-similar solutions to the Euler equations for the question of non-uniqueness. Homogeneous solutions to the Euler equations, like the power-law vortex $\Omega(\xi)=|\xi|^{-\alpha}$, are self-similar solutions to the Euler equations that are also stationary in the “physical coordinates”. Most of the qualitative properties of stationary homogeneous solutions to the two-dimensional Euler equations have been catalogued in the work [11] of Luo and Shvydkoy.
@@
The single spiral solution has been studied with mathematical rigor by several authors. Elling first proved in [9] the existence of self-similar solutions to the Euler equations from data that are small perturbations of the power-law vortex and $m$-fold rotationally symmetric with large $m$. His result was later refined by the authors of [7] and [13].
While the results in [9], [10], [13], and [7] are technically impressive, they leave the question of non-uniqueness (and even the question of how to approach this problem) completely open. In particular, the non-uniqueness scenario proposed in [4] and [5] requires the existence of a self-similar solution to the Euler equations whose pseudo-velocity has more than one stagnation point. On the other hand, all the previous constructions of self-similar spirals essentially depended on small perturbations of the inviscid power-law vortex. Consequently, all the solutions they construct have self-similar pseudo-velocity $U-\frac{\xi}{\alpha}$ with a single (spiral) stagnation point.
We believe it is worthwhile for the non-uniqueness program to discover self-similar profiles with pseudo-velocity having more than one stagnation point, and this is the goal of this note. We construct the first example of such a self-similar solution to the two-dimensional incompressible Euler equations.
@@
For any value of the scaling parameter $\alpha\in(0,1)$, there exists a self-similar profile $\Omega$ that is $(-\alpha)$-homogeneous and whose corresponding pseudo-velocity field $U-\frac{\xi}{\alpha}$ has two sink stagnation points away from the origin and a saddle stagnation point at the origin.
We call the self-similar profile we construct the two-sink solution. The pseudo-velocity field of the two-sink solution is pictured in Figure 1 and Figure 2. Keeping in mind the numerical experiments done in [4], we consider our two-sink solution as a plausible candidate for exhibiting the bifurcation behavior previously described. Our main theorem, in conjunction with the work showing the existence of self-similar solutions from generic data at infinity, demonstrates the feasibility of the non-uniqueness scenario proposed by Bressan and co-authors.
We note that our two-sink solution has velocity that is only Hölder continuous; however, this is not a peculiarity of our construction, but rather a consequence of the additional stagnation points. In particular, we prove in Proposition 3.7 below that any homogeneous self-similar solution to the two-dimensional incompressible Euler equations whose pseudo-velocity has more than one stagnation point has vorticity that is discontinuous along an infinite ray.
We thank Peter Constantin and Alexandru Ionescu for helpful comments on this note. We acknowledge Princeton University for its support. MPC also acknowledges the support of the National Science Foundation in the form of an NSF Graduate Research Fellowship.
@@ 2 Homogeneous Solutions of the Euler Equations and a Hamiltonian System
If $\lambda>0$, $P$ is a constant, and $\psi\in H^{1}(\mathbb{T})$ solves Equation (4), then $u=\nabla^{\perp}(r^{\lambda}\psi(\theta))\in L^{2}_{\text{loc}}$ is a steady state of the two-dimensional Euler equations. Since the solution $(u,p)$ is homogeneous, it is also a steady solution of the two-dimensional self-similar Euler equations (Equation (2)) with self-similar variable $\xi=x/t^{1/\alpha}$ where $\alpha=2-\lambda$. In general, we find homogeneous solutions $\Psi=r^{2-\alpha}\psi(\theta)$ by solving the ordinary differential equation given in Equation (4), which turns out to be a Hamiltonian system.
To construct the two-sink solution, we find a homogeneous solution with pseudo-velocity vector field $\nabla^{\perp}\Psi-\frac{1}{\alpha}\xi$ having two sink stagnation points. This is the first rigorous construction of solutions to the two-dimensional self-similar Euler equations whose pseudo-velocity has stagnation points other than the origin.
3 Construction of the Two-Sink Solution
In this section, we build upon the preliminary results on homogeneous solutions presented in the paper [11] of Luo and Shvdkoy, and we construct the two-sink solution. While Luo and Shvdkoy perform a complete classification of local solutions to the Hamiltonian system, their paper omits the construction of any particular non-trivial $2\pi$-periodic function $\psi(\theta)$. Our construction requires a fine asymptotic analysis of the period functions associated to the Hamiltonian system that will occupy us for the rest of this note.
3.1 Local solution
@@ 3.2 Gluing Local Solutions
When $1<\lambda<2$ and $P<0$, there are two distinct, positive, and maximal local solutions, which both connect $(0,\pm\sqrt{-2P})$ in the phase portrait of the Hamiltonian system. Let $\psi_{+}:[0,T_{+}]\to\mathbb{R}_{\geq 0}$ and $\psi_{-}:[0,T_{-}]\to\mathbb{R}_{\geq 0}$ be maximal local solutions corresponding to $B=1$ and B=1B=-1 respectively. Then, we can glue two solutions $\psi_{+}$ and $\psi_{-}$ at the endpoint to obtain a solution $\psi:[0,T_{+}+T_{-}]\to\mathbb{R}$ to the original differential equation in Equation (4). From the construction, $\psi\in C^{1}$ and since
@@
we in fact get $\psi\in C^{1,2-\frac{2}{\lambda}}$.
@@
Remark 3.2.
It is also possible to glue in the order $\psi_{+}$, $\psi_{+}$, $\psi_{-}$, $\psi_{-}$.
Remark 3.3.
If it happens that $T_{+}+T_{-}=\pi$, then $\psi$ is a $\pi$-periodic solution. Therefore, it gives a $2$-fold symmetric self-similar solution to the two-dimensional Euler equations.
Now our goal is to find values of $(\lambda,P)$ that make $T_{+}+T_{-}=\pi$. Since we have an analytic expression for the streamlines, $T_{\pm}$ can be written as an integral:
$T_{\pm}=\int_{0}^{1}\frac{2}{\sqrt{\lambda^{2}(1-s^{2})\mp x_{\pm}^{-2/\lambda}(1-s^{2-2/\lambda})}}\,ds,$
@@
For each $\lambda\in(1,2)$, there exists a $\pi$-periodic solution ($\psi$, $P$) of Equation (4) obtained by gluing $\psi_{+}$ and $\psi_{-}$.
@@
For each $\lambda\in(1,2)$, $x_{\pm}$ and $T_{\pm}$ are functions of $P\in(-\infty,0)$. As $P\to 0^{-}$, we have
$x_{+}^{-\frac{2}{\lambda}}(P=0)=\lambda^{2},\quad x_{-}(P=0)=0,$
and
$T_{+}(P=0)=\int_{0}^{1}\frac{2}{\lambda\sqrt{s^{2-2/\lambda}-s^{2}}}\,ds=\int_{0}^{1}\frac{2}{\sqrt{1-t^{2}}}\,dt=\pi,\quad T_{-}(P=0)=0.$
As $P\to$-$\infty$, we have $x_{\pm}(P=-\infty)=\infty$ and
$T_{\pm}(P=-\infty)=\frac{2}{\lambda}\int_{0}^{1}\frac{1}{\sqrt{1-s^{2}}}\,ds=\frac{\pi}{\lambda}.$
Thus, $T=T_{-}+T_{+}$ is a smooth function of $P\in(-\infty,0)$ such that
$T(P=0)=\pi,\quad T(P=-\infty)=\frac{2\pi}{\lambda}>\pi.$
In the following section, we show that $T$ decreases as P-P increases near $P=0$. Therefore, there exists $P_{*}<0$ such that $T(P=P_{*})=\pi$. ∎
We can now finish the proof of Theorem Theorem.
Theorem 3.5.
For each $\alpha\in(0,1)$, there exists a $\pi$-periodic solution $\Psi$ of Equation (2) whose corresponding pseudo-velocity $\nabla^{\perp}\Psi-\frac{\xi}{\alpha}$ has two sink stagnation points away from the origin and a saddle stagnation point at the origin.
Proof.
We have already shown in Theorem 3.4 that for every $\lambda\in(1,2)$, there exists a $\pi$-periodic solution ($\psi$, $P$) of Equation (4). The corresponding pseudo-velocity is:
@@
We can certainly choose some unique $r>0$ so that this is so. By the two fold reflection symmetry, there are exactly two stagnation points. Now, it is not difficult to see that the eigenvalues of $\nabla u$ at the two stagnation points away from the origin are precisely 1-1 and $\tfrac{-\lambda}{2-\lambda}$, so each of these stagnation points away from the origin is a sink. Lastly, the origin is a saddle since the flow is coming out (in) in the direction where $\psi^{\prime}<0~(>0)$. ∎
Remark 3.6.
Suppose $\Psi=r^{\lambda}\psi(\theta)$ with $\lambda\in(1,2)$ is a homogeneous steady state. The pseudo-velocity field $\nabla^{\perp}\Psi-\frac{\xi}{\alpha}$ has a stagnation point at each angle $\theta$ where $\psi(\theta)$ changes sign from positive to negative. Indeed, all stagnation points except the origin are equidistant from the origin since $|\psi^{\prime}(\theta)|=\sqrt{-2P}$ whenever $\psi(\theta)=0$.
@@
Let $\alpha\in(0,1)$. Suppose that $U(\xi)$ is a homogeneous self-similar solution to the two-dimensional Euler equations. Suppose in addition that the corresponding pseudo-velocity
$\tilde{U}(\xi):=U(\xi)-\frac{\xi}{2-\lambda}$
has a stagnation point in addition to the one at the origin $\xi=0$. Then the vorticity $\textrm{curl }\,U$ is discontinuous on a ray from the origin.
@@
As we observed in the proof of Theorem 3.5, if the pseudo-velocity $\tilde{U}$ has a stagnation point for $r>0$, the it necessarily occurs at an angle $\theta$ where the stream function $\psi(\theta)=0$ and $\psi^{\prime}(\theta)<0$. It is therefore not difficult to see that both the power-law vortex and the power-law shear flow are inadmissible (for one $\psi^{\prime}=0$ always, for the other $\psi^{\prime}=0$ whenever $\psi=0$). The only candidates remaining are obtained by gluing. As we previously observed, any gluing of local solutions yields at most $\psi\in C^{1,2-2/\lambda}$ regularity. In particular, the vorticity is discontinuous along the ray corresponding to the angle at which the stagnation points of $\tilde{U}$ occur.

### v2

# Multi-Sink Solutions to the Self-Similar Euler Equations

@@ Abstract
We construct examples and provide a classification of self-similar solutions to the two-dimensional incompressible Euler equations whose pseudo-velocity fields possess more than one stagnation point. These solutions are also homogeneous steady states of the Euler equations. In contrast, we prove that any homogeneous self-similar solution with bounded vorticity away from the origin necessarily admits only a single stagnation point, located at the origin. The solutions we construct develop velocity cusps along rays from the origin, and this allows for additional stagnation points of the pseudo-velocity field.
@@ 1 Introduction
$\left\{\begin{array}[]{r}\partial_{t}u+(u\cdot\nabla)u+\nabla p=0,\\ u(t=0)=u_{0}.\end{array}\right$. (1)
One solves for the velocity vector field $u(x,t)$ and scalar pressure field $p(x,t)$ starting from initial data $u_{0}(x)$. The incompressible Euler equations in two-dimensions model the motion of inviscid fluids in thin plates. Due to a transport structure of the equations, the vorticity, $\textrm{curl }u$, is transported by the velocity $u$, which facilitates the existence theory for weak solutions starting from quite irregular initial data (for example, the existence for $L^{1}\cap L^{p}$ vorticity was proven in [8]). On the other hand, uniqueness of solutions is only known when the vorticity is bounded, and non-uniqueness is only known (by convex integration arguments) due to a lack of energy conservation for $C^{1/3^{-}}$ velocity. More recently, the authors of [6] used convex integration to prove non-uniqueness for $L^{1+}$ vorticity data on $\mathbb{T}^{2}$, but these are not solutions in the usual distributional sense. The question whether unbounded and integrable ($L^{1}\cap L^{p}$) vorticity data guarantees the uniqueness of weak solutions is open and referred to as the “Yudovich Problem”.
@@
$\left\{\begin{array}[]{r}(\nabla^{\perp}\Psi-\frac{1}{\alpha}\xi)\cdot\nabla_{\xi}\Omega=\Omega,\\ \Delta\Psi=\Omega.\end{array}\right$. (2)
@@
$\omega(x,t):=t^{-1}\Omega(xt^{-1/\alpha}),$
@@
Bifurcation in the self-similar variables is believed to be a plausible non-uniqueness mechanism for the Euler equations, especially since Vishik used a similar scenario to prove the non-uniqueness of solutions with time-integrable forcing in [15] and [16] (a result further refined by the authors of [1]). We now describe the bifurcation behavior necessary for non-unique solutions to arise. If one can construct two different self-similar profiles with the same “data at infinity” in the self-similar variables, then one has proven the existence of two different self-similar solutions to the Euler equations from the same homogeneous data.
@@
We have argued the importance of self-similar solutions to the Euler equations for the question of non-uniqueness. Homogeneous solutions to the Euler equations, like the power-law vortex with vorticity $\Omega(\xi)=|\xi|^{-\alpha}$, are self-similar solutions to the Euler equations that are also stationary in the “physical coordinates”. Most of the qualitative properties of stationary homogeneous solutions to the two-dimensional Euler equations have been catalogued in the work [12] of Luo and Shvydkoy.
@@
The single spiral solution has been studied with mathematical rigor by several authors. Elling first proved in [10] the existence of self-similar solutions to the Euler equations from data that are small perturbations of the power-law vortex and $m$-fold rotationally symmetric with large $m$. His result was later refined by the authors of [7] and [14]. While the results in [10], [11], [14], and [7] are technically impressive, they leave the question of how to approach the Yudovich problem completely open. In particular, the non-uniqueness scenario proposed in [4] and [5] requires the existence of a self-similar solution to the Euler equations whose pseudo-velocity has multiple stagnation points. On the other hand, all the previous constructions of self-similar spirals essentially depended on small perturbations of the inviscid power-law vortex. Consequently, all the solutions they construct have self-similar pseudo-velocity $U-\frac{\xi}{\alpha}$ with a single spiral stagnation point.
We believe it is worthwhile for the non-uniqueness program to discover self-similar profiles whose pseudo-velocities have more than one stagnation point. To this end, we construct and classify homogeneous self-similar profiles with pseudo-velocity fields having multiple sinks away from the origin. We call these self-similar profiles multi-sink solutions. As a first example of a multi-sink solution, we have the following consequence of Theorem 3.3 and Theorem 3.5 that we prove below.
@@
For any scaling parameter $\alpha\in(0,1)$, there exist $(-\alpha)$-homogeneous self-similar vorticity profiles $\Omega$ for which the corresponding pseudo-velocity field $U-\frac{\xi}{\alpha}$ has multiple sink stagnation points away from the origin and a saddle point at the origin. In particular, up to isometries of the plane, there exists a unique such profile that has exactly two sinks and is $2$-fold rotationally symmetric.
The construction of multi-sink solutions to the incompressible two-dimensional Euler equations is based on gluing together solutions that are locally defined on sectors. There are many ways to perform such gluings with the hope of constructing multi-sink solutions. In the case of $\alpha\in[1/2,1)$, we in fact provide a complete classification of multi-sink solutions. Taken together with the classification done in [12] of homogeneous steady states that are not obtained by gluing, our Theorem 3.3, completes the classification of all $(-\alpha)$-homogeneous steady states of the two-dimensional Euler equations when $\alpha\in[1/2,1)$. In the case when $\alpha\in(0,1/2)$, we are also able to construct many examples of multi-sink solutions, and a finer analysis of certain “transcendental elliptic integrals” used in the construction would probably lead to a complete classification in this case as well.
Our multi-sink solutions have velocity that is only Hölder continuous, due to the gluing procedure; however, this is not a peculiarity of our construction, but rather a consequence of the additional stagnation points. In particular, we prove in Proposition 3.6 below that any homogeneous self-similar solution to the two-dimensional incompressible Euler equations whose pseudo-velocity has more than one stagnation point has vorticity that is discontinuous and unbounded along an infinite ray.
We particularly note the existence of two distinct multi-sink solutions that are $2$-fold rotationally symmetric and possess two sinks away from the origin. The first example, which is additionally $4$-fold rotationally symmetric, was utilized as a barrier function in a technical argument by the authors of [9]. The second example is the one highlighted in the last sentence of our Theorem, and we refer to this as the two-sink solution. This solution is of special interest since it appears to converge to the power-law shear flow with velocity $(y|y|^{-\alpha},0)$ at a quadratic rate as $\alpha\to 0^{+}$. In this sense, the two-sink solution can be viewed as bifurcating from the power-law shear flow. We provide a combination of numerical and rigorous analytical evidence for this convergence in Section 4.
The pseudo-velocity field of the two-sink solution is pictured in Figure 1 and Figure 2. Keeping in mind the numerical experiments done in [4], we consider our two-sink solution as a plausible candidate for a non-uniqueness mechanism. The theorem we state above, in conjunction with the work showing the existence of self-similar solutions from generic data at infinity, demonstrates the feasibility of the non-uniqueness scenario proposed by Bressan and co-authors.
Acknowledgments. We thank Peter Constantin and Alexandru D. Ionescu for helpful comments on this note. We thank Tarek M. Elgindi and Yupei Huang for pointing out their work in [9]. We acknowledge Princeton University for its support. MPC also acknowledges the support of the National Science Foundation in the form of an NSF Graduate Research Fellowship.
@@ 2 Homogeneous Solutions of the Euler Equations and a Hamiltonian System
If $\lambda>0$, $P$ is a constant, and $\psi\in H^{1}(\mathbb{T})$ solves Equation (3), then $u=\nabla^{\perp}(r^{\lambda}\psi(\theta))\in L^{2}_{\text{loc}}$ is a steady state of the two-dimensional Euler equations. Since the solution $(u,p)$ is homogeneous, it is also a steady solution of the two-dimensional self-similar Euler equations (Equation (2)) with self-similar variable $\xi=x/t^{1/\alpha}$ where $\alpha=2-\lambda$.
In general, we find homogeneous solutions $\Psi=r^{2-\alpha}\psi(\theta)$ by solving the ordinary differential equation given in Equation (3), which turns out to be a Hamiltonian system. This method for constructing homogeneous solutions to the two-dimensional Euler equations was used in [12] and is probably known to earlier mathematicians as well. To construct multi-sink solutions, we find homogeneous solutions with pseudo-velocity vector field $\nabla^{\perp}\Psi-\frac{1}{\alpha}\xi$ having several sink stagnation points.
3 Construction of Multi-Sink Solutions
In this section, we build upon the preliminary results on homogeneous solutions presented in the paper [12] of Luo and Shvdkoy, and we construct multi-sink solutions. While Luo and Shvdkoy perform a complete classification of local solutions to the Hamiltonian system, their paper omits the construction of any particular non-trivial $2\pi$-periodic function $\psi(\theta)$. Our construction requires a fine asymptotic analysis of the period functions associated to the Hamiltonian system that will occupy us for the rest of this note.
3.1 Local Solutions
@@ 3.2 Gluing Local Solutions
Let $\psi_{1}:[0,T_{1}]\to\mathbb{R}_{\geq 0}$ and $\psi_{2}:[0,T_{2}]\to\mathbb{R}_{\geq 0}$ be maximal local solutions with the same value of $P<0$. Then, we can glue two solutions $\psi_{1}$ and $\psi_{2}$ at the endpoint to obtain a solution $\psi:[0,T_{1}+T_{2}]\to\mathbb{R}$ to the original differential equation in Equation (3). From the construction, $\psi\in C^{1}$ and since
@@
we in fact get $\psi\in C^{1,2-\frac{2}{\lambda}}$. In this way, we can glue any even number of maximally defined local solutions to obtain a $C^{1,2-\frac{2}{\lambda}}$ solution to Equation (3). When $1<\lambda<2$ and $P<0$, there are two distinct, positive, and maximal local solutions, which both connect $(0,\pm\sqrt{-2P})$ in the phase portrait of the Hamiltonian system. We denote by $\psi_{\pm}$, the maximal local solution corresponding to $B=\pm 1$.
@@
Let $T_{\pm}$ be the lifespans of maximal local solutions corresponding to $B=\pm 1$. Since we have an analytic expression for the streamlines, $T_{\pm}$ can be written as an integral:
$T_{\pm}=\int_{0}^{1}\frac{2}{\sqrt{\lambda^{2}(1-s^{2})\mp x_{\pm}^{-2/\lambda}(1-s^{2-2/\lambda})}}\,ds,$ (9)
@@
The integrals $T_{\pm}$ are the “transcendental elliptic integrals” we referred to in the introduction.
Lemma 3.2.
If $1<\lambda<2$, then the integrals $T_{\pm}$ are continuous monotonic functions of $P\in(-\infty,0]$, where $T_{+}(P=0)=\pi$, $T_{-}(P=0)=0$, and $\lim_{P\to-\infty}T_{\pm}=\frac{\pi}{\lambda}$.
Proof.
Since $\lambda^{2}=x_{+}^{-\frac{2}{\lambda}}-2Px_{+}^{-2}$, we observe that $x_{+}^{-\frac{2}{\lambda}}$ as a function of $P\in[0,\infty)$ is continuous, monotonically decreasing, and
$x_{+}^{-\frac{2}{\lambda}}(P=0)=\lambda^{2},\quad x_{+}^{-\frac{2}{\lambda}}(P=-\infty)=0.$
When $x_{+}^{-\frac{2}{\lambda}}\in[\lambda^{2},\infty]$, the integral expression of $T_{+}$ in Equation (9) is integrable, so $T_{+}$ is continuous and monotonically decreasing in $P\in[0,\infty)$. In particular, the values of $T_{+}$ at the endpoints $P=0$ and $P=$-$\infty$ are evaluated to be
$T_{+}(P=0)=\int_{0}^{1}\frac{2}{\lambda\sqrt{s^{2-2/\lambda}-s^{2}}}\,ds=\int_{0}^{1}\frac{2}{\sqrt{1-t^{2}}}\,dt=\pi,$
and
$T_{+}(P=-\infty)=\frac{2}{\lambda}\int_{0}^{1}\frac{1}{\sqrt{1-s^{2}}}\,ds=\frac{\pi}{\lambda}.$
Similarly, $x_{-}$ as a function of $P\in[0,\infty)$ is continuous, monotonically increasing, and
$x_{-}(P=0)=0,\quad x_{-}(P=-\infty)=\infty.$
Thus, $T_{+}$ is continuous and monotonically increasing in $P\in[0,\infty)$. Moreover, the values of $T_{+}$ at the endpoints $P=0$ and $P=$-$\infty$ are
$T_{-}(P=0)=x_{-}^{1/\lambda}(P=0)\int_{0}^{1}\frac{2}{\lambda}\frac{ds}{\sqrt{1-s^{2}}}=0,$
and
$T_{-}(P=-\infty)=\frac{2}{\lambda}\int_{0}^{1}\frac{1}{\sqrt{1-s^{2}}}\,ds=\frac{\pi}{\lambda}.$
@@
If $1<\lambda<2$, then a $2\pi$-periodic solution to Equation (3) can be obtained by gluing any of the following combinations in any order with alternating signs:
  1. (i)
    $2k$ copies of $\psi_{-}$, for $k\geq 2$.
  2. (ii)
    One copy of $\psi_{+}$ and 2k12k-1 copies of $\psi_{-}$, for $k\geq 2$.
  3. (iii)
    Two copies of $\psi_{+}$.
  4. (iv)
    Two copies of $\psi_{+}$ and two copies of $\psi_{-}$.
In particular, these possibilities together with (v) below, constitutes the complete list of possible gluing combinations for $1<\lambda\leq\frac{3}{2}$.
  1. (v)
    Two copies of $\psi_{+}$ and $2k$ copies of $\psi_{-}$, for $k\geq 2$.
Remark 3.4.
(a) In the case (iii), we necessarily have that $P=0$ since $T_{+}(P=0)=\pi$ and $T_{+}$ is otherwise smaller than $\pi$. The $P=0$ solution to Equation (3) is the homogeneous shear flow solution, either $\Psi=|y|^{\lambda}$ or $y|y|^{\lambda-1}$.
(b) The case (iv), if we glue in the order $\psi_{+},-\psi_{-},\psi_{+},-\psi_{-}$, yields a $2$-fold symmetric homogeneous solution to the two-dimensional Euler equations, which we refer to as the two-sink solution.
(c) In the case (i) with $k=2$, the resulting solution gives an exactly $4$-fold rotationally symmetric homogeneous steady solution to the two-dimensional Euler equations. This case was observed by Elgindi and Huang in the appendix of [9].
Proof of Theorem 3.3. We will show in every case that there exists $P<0$ such that sum of the lifespans equal $2\pi$.
Case (i). Since $T_{-}(P=0)=0$ and $T_{-}(P=-\infty)=\frac{\pi}{\lambda}>\frac{\pi}{2}$, there exists $P<0$ such that $T_{-}(P)=\frac{\pi}{k}$ for any $k\geq 2$.
Case (ii). Notice that $T_{+}(P=0)+(2k-1)T_{-}(P=0)=\pi$ and $T_{+}(P=-\infty)+(2k-1)T_{-}(P=0)=\frac{2k\pi}{\lambda}>2\pi$. Thus, there exists $P<0$ such that $T_{+}(P=0)+(2k-1)T_{-}(P=0)=2\pi$ for any $k\geq 2$.
Case (iii). Simply $T_{+}(P=0)=\pi$.
Case (iv). Notice that $T_{+}(P=0)+T_{-}(P=0)=\pi$ and $T_{+}(P=-\infty)+T_{-}(P=-\infty)=\frac{2\pi}{\lambda}>\pi$. We prove in the next section that $T_{+}+T_{-}$ is a decreasing function of P-P near $P=0$. This shows that there exists $P<0$ such that $T_{+}(P=0)+T_{-}(P=0)=\pi$.
When $1<\lambda\leq\frac{3}{2}$. We see that $T_{+}>\frac{\pi}{\lambda}\geq\frac{2\pi}{3}$. Thus, there cannot be more than two copies $\psi_{+}$ if we are to obtain a $2\pi$-periodic solution. In addition, since $T_{\pm}<\pi$ for $P<0$, it is impossible to have only two copies of $\psi_{-}$, nor is it possible to have one copy of $\psi_{+}$ and $\psi_{-}$ each. It follows that the cases (i), (ii), (iii), (iv), and (v) are only possible combinations. Indeed, we prove in the next section that $T_{+}+kT_{-}$ is a decreasing function of P-P near $P=0$ when $1<\lambda<\frac{3}{2}$, so the case (v) is valid. ∎
We can now finish the proof of the theorem we stated in the introduction.
Theorem 3.5.
Suppose $P<0$, and let $\psi$ be a $2\pi$-periodic solution to Equation (3) obtained by gluing $2k$ maximal local solutions. Consider the corresponding homogeneous steady state $\Psi=r^{\lambda}\psi(\theta)$. Then, the pseudo-velocity field $\nabla^{\perp}\Psi-\frac{\xi}{\alpha}$ has a stagnation point at each angle $\theta$ where $\psi(\theta)$ changes sign from positive to negative. Moreover, all $k$ stagnation points other than the origin are sinks located at a distance of $\alpha^{\frac{1}{\alpha}}(-2P)^{\frac{1}{2\alpha}}$ from the origin. Finally, the origin is a saddle point.
@@
The corresponding pseudo-velocity is:
@@
Therefore, on the angle $\theta$ with $\psi(\theta)=0$, $\psi^{\prime}(\theta)<0$, there is a stagnation point with distance $r$ determined by the equation above. Since $|\psi^{\prime}(\theta)|=\sqrt{-2P}$ given that $\psi(\theta)=0$, we show that the distance $r=\alpha^{\frac{1}{\alpha}}(-2P)^{\frac{1}{2\alpha}}$.
Now, it is not difficult to see that the eigenvalues of $\nabla(U-\xi/(2-\lambda))$ at the two stagnation points away from the origin are precisely 1-1 and $\tfrac{-\lambda}{2-\lambda}$, so each of these stagnation points away from the origin is a sink. Lastly, the origin is a saddle since the flow is coming out (in) in the direction where $\psi^{\prime}<0~(>0)$. ∎
@@
Let $\alpha\in(0,1)$. Suppose that $U(\xi)$ is a $(1-\alpha)$-homogeneous steady solution to the two-dimensional Euler equations. Suppose in addition that the corresponding pseudo-velocity
$\tilde{U}(\xi):=U(\xi)-\frac{\xi}{\alpha}$
has a stagnation point in addition to the one at the origin $\xi=0$. Then the vorticity $\textrm{curl}~U$ is discontinuous on a ray from the origin.
@@
As we observed in the proof of Theorem 3.5, if the pseudo-velocity $\tilde{U}$ has a stagnation point for $r>0$, the it necessarily occurs at an angle $\theta$ where the stream function $\psi(\theta)=0$ and $\psi^{\prime}(\theta)<0$. It is therefore not difficult to see that both the power-law vortex and the power-law shear flow are inadmissible (for one $\psi^{\prime}=0$ always, for the other $\psi^{\prime}=0$ whenever $\psi=0$). The only candidates remaining are obtained by gluing. As we previously observed, any gluing of local solutions yields at most $\psi\in C^{1,2-2/\lambda}$ regularity. In particular, the vorticity is discontinuous along the ray corresponding to the angle at which the stagnation points of $\tilde{U}$ occur. ∎